OpenStudy (anonymous):

Solve the equation for x: 6x + 12y = 5. x = 5 – 12y x = 5/6 – 12y x = 5/6 – 2y x = 5 – 2y can someone explain me how this works?

OpenStudy (anonymous):

okk

OpenStudy (anonymous):

x = 5/6 – 2y

OpenStudy (anonymous):

6x + 12y = 5. 6x = 5-12y x = (5-12y)/6 = 5/6 - 12y/6 = 5/6 - 2y

OpenStudy (anonymous):

thanks man

OpenStudy (anonymous):

u r welcome ☺

OpenStudy (anonymous):

can you also help me with this plz it sais solve for P 8z-p/5=3z p = 5 – 11z p = 5z - 5 p = – 5z p = -7z

OpenStudy (anonymous):

8z - p/5 = 3z 8z - 3z = p/5

OpenStudy (anonymous):

8z-p/5=3z 8z-3z=p/5 5z = p/5 5z*5 = p 25z = p

OpenStudy (anonymous):

soo the answer is the 3d option p=-5?

OpenStudy (anonymous):

ans is p=25z

OpenStudy (anonymous):

they dont give me that option :/ only have the 4 above sorry man i think you went wrong somewhere

OpenStudy (anonymous):

8z - p/5 = 3z 8z - 3z = p/5 5z = p/5 z = p/25 wrong options

OpenStudy (anonymous):

p = 25z